# Early Risk Prediction of Pediatric Cardiac Arrest from Electronic Health Records via Multimodal Fused Transformer

### v1

@@ I Introduction
To address the above mentioned challenges, we propose our innovative approach, namely PEDiatric Cardiac Arrest prediction via Fused Transformer (PedCA-FT), based on a novel tabular-textual multimodal fusion strategy [14, 15], and the powerful Transformer backbone model for each modality. We view the heterogeneous and multi-resolution EHR risk factors for pediatric CA as two complementary modalities, tabular and textual, each characterized by its own intrinsic data structure. Specifically, we employ a dedicated tabular Transformer [16, 17] to effectively handle high-dimensional static and aggregated longitudinal tabular features, while a pre-trained textual EHR Transformer [18] processes the textual representations derived from the original EHR data. A fusion Transformer is then used to integrate the modality-specific representations, ultimately computing the final probability of CA onset risk. We evaluate the effectiveness of our PedCA-FT on a curated cohort from the CHOA-CICU database comprising 3,566 pediatric patients with a $4.0\%$ incidence of CA. Our model, along with ten other AI models, is compared using 5-fold cross-validation. Through this comprehensive evaluation, our proposed approach marginally outperforms all compared models. Furthermore, a feature importance analysis reveals that several of the identified risk factors align well with clinical knowledge for pediatric CA, as confirmed by our clinical collaborators.
@@ II-B Data Source and Cohort Selection
CHOA-CICU database is a private pediatric intensive care database extracted from electronic health record for all pediatric patients (age less than 18 years) admitted from 1/1/2018 to 12/31/2023 to a large, quaternary, academic Pediatric Cardiac Intensive Care Unit (CICU) at Children’s Hospital at Atlanta and Emory University. We obtain 9 static risk factors from CHOA EHR, including: “gender, ethnicity, first race, second race, primary language, respiration distress, age, admission diagnoses, admission ICD 19 codes”. Among 184 temporal risk factors, we obtain 129 vital signs and nursing assessments from flow sheet after removing non-relevant items and removing missing rate $>0.5$ items. We also obtain 49 lab test results after removing missing rate $>0.5$ items, and 5 medications after removing missing rate $>0.7$ items. Table I briefly presents the number of patients included in the two data sources. We further conduct cohort selection to exclude certain patients (exactly the same as submitted abstract).
@@ III-A Tabular Transformer
where $\phi:\mathbb{R}^{d_{t}\times\tau}\rightarrow\mathbb{R}^{d_{t}}$ denotes the aggregation operator that maps the time-series features into tabular features. A wide range of aggregation operation can be used for $\phi$, and we opt to $LAST(\cdot)$ which keeps the last observed element from the whole time-series. The input of tabular-transformer $\mathcal{T}_{tab}$ is $\mathbf{x}_{t\leq\tau}=(\mathbf{x}^{(s)},\mathbf{x}^{(t)}_{\text{agg}})$ after applying the aggregation using Eq. (1).

### v2

@@ I Introduction
To address the above mentioned challenges, we propose our innovative approach, namely PEDiatric Cardiac Arrest prediction via Fused Transformer (PedCA-FT), based on a novel tabular-textual multimodal fusion strategy [14, 15], and the powerful Transformer backbone model for each modality. We view the heterogeneous and multi-resolution risk factors as two complementary modalities, tabular and textual, each characterized by its own intrinsic data structure. Specifically, we employ a dedicated tabular Transformer [16, 17] to effectively handle high-dimensional static and aggregated longitudinal tabular features, while a pre-trained textual EHR Transformer [18] processes the textual representations [19] derived from the original EHR data. A fusion Transformer is then used to integrate the modality-specific representations, ultimately computing the probability of CA onset risk. We evaluate the effectiveness of our PedCA-FT on a curated cohort from the CHOA-CICU database comprising 3,566 pediatric patients with a $4.0\%$ incidence of CA. Our model, along with ten other AI models, is compared using 5-fold cross-validation. Through this comprehensive evaluation, our proposed approach marginally outperforms all compared models. Furthermore, a feature importance analysis reveals that several of the identified risk factors align well with clinical knowledge for pediatric CA, as confirmed by our clinical collaborators.
@@ II-B Data Source and Cohort Selection
CHOA-CICU database is a private pediatric intensive care database extracted from electronic health record for all pediatric patients (age less than 18 years) admitted from 1/1/2018 to 12/31/2023 to a large, quaternary, academic Pediatric Cardiac Intensive Care Unit (CICU) at Children’s Hospital at Atlanta and Emory University. We obtain 9 static risk factors from CHOA EHR, including: “gender, ethnicity, first race, second race, primary language, respiration distress, age, admission diagnoses, admission ICD 19 codes”. Among 184 temporal risk factors, we obtain 129 vital signs and nursing assessments from flow sheet after removing non-relevant items and removing missing rate $>0.5$ items. We also obtain 49 lab test results after removing missing rate $>0.5$ items, and 5 medications after removing missing rate $>0.7$ items. Table I briefly presents the number of patients included in the two data sources.
@@ III-A Tabular Transformer
where $\phi:\mathbb{R}^{d_{t}\times\tau}\rightarrow\mathbb{R}^{d_{t}}$ denotes the aggregation operator that maps the time-series features into tabular features. A wide range of aggregation operation can be used for $\phi$, and we opt to $LAST(\cdot)$ which keeps the last observed element from the whole time-series. The input of tabular-transformer $\mathcal{T}_{tab}$ is $\mathbf{x}_{t\leq\tau}=(\mathbf{x}^{(s)},\mathbf{x}^{(t)}_{\text{agg}})$ after applying Eq. (1).